OpenStudy (anonymous):

Solve this quadratic equation using the quadratic formula. x2 + 8x - 5 = 0

OpenStudy (anonymous):

What do you think ??

OpenStudy (anonymous):

Have you tried it once ??

OpenStudy (anonymous):

i dont know, no.

OpenStudy (anonymous):

Here: a = 1, b = 8 and c = -5 Firstly find Discriminant using: \[D = b^2 - 4ac\] Then find x by using: \[x = \frac{-b \pm \sqrt{D}}{2a}\]

OpenStudy (anonymous):

i have no clue

OpenStudy (anonymous):

Find D: \[D = b^2 - 4ac\] Here : b = 8, a = 1 and c = -5.. Just plug in the values..

OpenStudy (swissgirl):

well water in eyes divided into 2 steps which is the easiest. Lets first fine out what our D equals

OpenStudy (swissgirl):

Like he said b=8 our a=1 and c=-5 \(8^2-4(1)(-5) =?\)

OpenStudy (swissgirl):

Put this into a calc and see what you get

OpenStudy (swissgirl):

Are you there?

OpenStudy (anonymous):

i solved it. thanks

OpenStudy (swissgirl):

Great :D